A tank holds 800 gallons of water. The gauge indicates that the tank is ¾ full; then 100 gallons are drained from the tank to fill a pool 7/8 full. Next, 1/3 of a tank of water is added back in. What fraction of the tank is filled with water?
I have figured out that 800 gallons = 3/4 full (tank) and that 3/4 = 6/8. I do not know how to solve the problem from here!

Found 2 solutions by Theo, stanbon:
Answer by Theo(2973)

A tank holds 800 gallons of water. The gauge indicates that the tank is ¾ full; then 100 gallons are drained from the tank to fill a pool 7/8 full. Next, 1/3 of a tank of water is added back in. What fraction of the tank is filled with water?

Let T represent a full tank = 800 gallons of water.
Let P represent a full pool = ???

3/4 * T - 100 = 7/8 * P

This formula says that three quarters of the tank minus 100 gallons equals 7/8 of the capacity of the pool.

You are adding 1/3 of a tank of water back in to the tank ?????

This problem apparently doesn't have anything to do with the pool.

Here's why I think that way.

Your tank is 3/4 full meaning it have 800*3/4 = 600 gallons in it.

you drain 100 gallons from it so you are left with 500 gallons in the tank.

you add 1/3 of a tank full of water back into it, so you are adding 800/3 = 266 and 2/3 gallons into it.

500 + 266 and 2/3 = 766 and 2/3 gallons / 800 = .9583333333 percent full.

in fractions, you would have (2300/3) / (2400/3) which is the same as 2300/2400.

2300/2400 is the same as 23/24 = .95833333333 so your fraction of the tank that is full would be 23/24.

The pool had nothing to do with it other than to add extraneous information as far as I could see.










Answer by stanbon(48558)

A tank holds 800 gallons of water.
The gauge indicates that the tank is ¾ full = 600 gallons
; then 100 gallons are drained from the tank = 500
to fill a pool 7/8 full.
Next, 1/3 of a tank of water is added back in = 500 + (1/3)800 = 766 2/3 gallons
--------
What fraction of the tank is filled with water?
-----
Fraction = (766 2/3)/800 = [2300/3]/800 = 2300/2400 = 23/24
==================================================================
Cheers,
Stan H.